# Six-Point Method for Multi-Camera Systems with Reduced Solution Space

### v1

@@ 8.1 Numerical Stability
Figure 9: Rotation and translation error with varying image noise for the inter-camera case. We designed a simulated two-camera rig composed of two perspective cameras. The orientations of the two perspective cameras are roughly forward-facing with random perturbation. The three rows correspond to forward, sideways, and random motions, respectively.
@@
Figure 10: Rotation and translation error with varying image noise for the intra-camera case. We designed a simulated two-camera rig composed of two perspective cameras. The orientations of the two perspective cameras are roughly forward-facing with random perturbation. The three rows correspond to forward, sideways, and random motions, respectively.
8.2 Expected Runtime vs. Inlier Ratio Accounting for Numerical Stability
The runtime of RANSAC depends on the runtime of the minimal solver and iteration number. For the RANSAC methods, the iteration number is determined by the probability of choosing at least one inlier set. This calculation implicitly assumes that an accurate solution can be obtained given an inlier set. However, this is not always true when considering the numerical stability of the minimal solvers.
$(1-(1-\epsilon)^{s})^{N}=1-p,$ (46)
and
$N=\frac{\log(1-p)}{\log(1-(1-\epsilon)^{s})}.$ (47)
$(1-(p_{2}(1-\epsilon))^{s})^{\hat{N}}=1-p,$ (48)
and
$\hat{N}=\frac{\log(1-p)}{\log(1-(p_{2}(1-\epsilon))^{s})}.$ (49)
The left question is how to determine $p_{2}$ for a given minimal solver. For each minimal solver, we can empirically obtain the probability that the estimation errors $\varepsilon_{\mathbf{R}}$ and $\varepsilon_{\mathbf{t}}$ on noise-free observations are below specified thresholds. In this paper, we set the thresholds as $10^{-3}$. For the solvers 17pt-Li [35], 8pt-Kneip [24], 6pt-Stewénius [46], 6pt-Our-inter56, 6pt-Our-inter48 and 6pt-Our-intra, $p_{2}$ can be computed as 1.00, 0.09, 0.59, 0.99, 0.90 and 0.81, respectively. Fig. 8 shows RANSAC iteration numbers with respect to outlier ratio for success probability $99\%$.
8.3 Experiments on Synthetic Data
Following the experiments about numerical stability, two simulated scenarios are designed and tested for the synthetic experiments, including scenarios for a two-camera rig and a generalized camera. The proposed solvers are compared with state-of-the-art solvers including 17pt-Li [35], 8pt-Kneip [24], and 6pt-Stewénius [46].
8.3.1 Synthetic Experiments for Two-camera Rigs
In this scenario, we design a simulated two-camera rig composed of two perspective cameras. The orientations of the two perspective cameras are roughly forward-facing with random perturbation. This setting is practical for autonomous driving with two front-facing cameras [10]. The performance of the solvers are tested in inter-camera and intra-camera configurations. The baseline length between the two simulated cameras is set to $1$ meter. The multi-camera reference frame is defined at the middle of the camera rig, and the translation between two multi-camera reference frames is $3$ meters. The resolution of the cameras is $640\times 480$ pixels, and the focal lengths are $400$ pixels. The principal points are set to the image center. The synthetic scene is made up of a ground plane and $50$ random planes. These planes are randomly generated in a cube of $[-5,5]\times[-5,5]\times[10,20]$ meters, which are expressed in the respective axis of the multi-camera reference frame. We randomly choose $50$ PCs from the ground plane and a PC from each random plane. Thus, there are $100$ PCs generated randomly in the synthetic data.
A total of $1000$ trials are carried out in the synthetic experiment. In each trial, $100$ PCs are generated randomly. Correspondences for the different methods are selected randomly. The error is measured on the best relative pose, which produces the most inliers within the RANSAC scheme. The RANSAC scheme also allows us to select the best candidate from multiple solutions. The median of errors is used to assess the rotation and translation errors. In this set of experiments, the translation direction between two multi-camera references is chosen to produce forward, sideways, or random motions. For each motion, the second view is perturbed by a random rotation. This random rotation is rotated around three axes in order, and the rotation angles range from $-10^{\circ}$ to $10^{\circ}$.
@@ 8.3.1 Synthetic Experiments for Two-camera Rigs
Fig. 9 demonstrates the performance of different solvers against image noise for the inter-camera case. A good method should have small median and variance for ${\varepsilon_{\mathbf{R}}}$, ${\varepsilon_{\mathbf{t}}}$, and $\varepsilon_{\mathbf{t},\text{dir}}$. We have the following observations. (1) The proposed 6pt-Our-inter solver provides better results than the comparative solvers. (2) The iterative optimization in 8pt-Kneip is susceptible to falling into local minima. It performs well for forward and random motions. However, it does not perform well for the sideways motion. (3) The linear solver 17pt-Li with fewer calculations have less round-off error than the proposed 6pt-Our-inter solver in noise-free cases. However, our method has better accuracy than the linear solver with the influence of image noise. This is also consistent with real-world data experiments.
@@
Fig. 10 demonstrates the performance of different solvers against image noise for the intra-camera case. We have the following observations. (1) Compared to the results in Fig. 9, the solvers using intra-camera correspondences generally perform worse than inter-camera correspondences, especially in recovering the metric scale of translation. (2) The proposed 6pt-Our-intra solver has better performance than the comparative solvers. (3) The 8pt-Kneip performs well in the forward motion of the multi-camera systems but poorly in the sideways and random motions.
8.3.2 Synthetic Experiments for Generalized Camera
@@
Table 9: The correspondence selection for different solvers. The pair $(i,i^{\prime})$ represents a PC that appears in the $i$-th camera in the first view and the $i^{\prime}$-th camera in the second view. The multiplier after a pair means the number of this match type.
@@
In this scenario, we design a simulated generalized camera composed of 12 omnidirectional cameras. The extrinsic parameters, including orientation and position, are totally random. Some of the comparison solvers use more PCs than the proposed solvers. The strategy of correspondence selection has a significant influence on their performance. We design a rule to select matches for different cases to guarantee fairness. Take the generic-camera configuration as an example. The match types are $(1,2)$, $(3,4)$, $(5,6)$, $(7,8)$, $(9,10)$, and $(11,12)$ for the proposed generic solver. Recall that pair $(i,i^{\prime})$ represents a PC that appears in the $i$-th camera in the first view and the $i^{\prime}$-th camera in the second view. Solvers 17pt-Li, 8pt-Kneip, and 6pt-Stewénius cyclically build matches according to the order of $(1,2)$, $(3,4)$, $(5,6)$, $(7,8)$, $(9,10)$, and $(11,12)$. Tab. 9 shows the correspondence selection for different solvers.
8.4 Experiments on KITTI Dataset
The empirical cumulative error distributions for KITTI sequence 00 are shown in Fig. 11. These values are calculated from the same values which were used for creating Table 3 in the paper. It can be seen that the proposed 6pt-Our-intra solver offers the best overall performance in comparison to state-of-the-art solvers.
@@ 8.4 Experiments on KITTI Dataset
8.5 Experiments on nuScenes Dataset
The performance of the solvers is also tested on the nuScenes dataset [6], which consists of consecutive keyframes from 6 cameras. This multi-camera system provides a full 360-degree field of view. We utilize all the keyframes of Part 1 for the evaluation, and there are 3376 images in total. The ground truth is given by a lidar map-based localization scheme. Similar to the experiments on KITTI dataset, the SIFT detector [38] is used to establish the PCs between consecutive views in six cameras. All the solvers are integrated into RANSAC to remove outlier matches of the feature correspondences.
Table 10: Rotation and translation error on nuScenes sequences (unit: degree).
@@ 8.5 Experiments on nuScenes Dataset
01 0.183  2.826 0.175   2.732 0.214   2.978 0.155  2.519
Tab. 10 shows the rotation and translation error of the proposed 6pt-Our-intra for the Part1 of nuScenes dataset. The median error is used to evaluate the estimation accuracy. It is demonstrated that the proposed 6pt-Our-intra offers the best performance among all the methods. In comparison with experiments on KITTI dataset, this experiment also demonstrates that 6pt-Our-intra can be directly used for the relative pose estimation for the systems with more cameras.
8.6 Experiments on EuRoC Dataset
We further validate the proposed solvers in an unmanned aerial vehicle environment. The EuRoC MAV dataset [4] is used to evaluate the performance of 6DOF relative pose estimation, which is recorded using a stereo camera mounted on a micro aerial vehicle. The sequences labeled from MH01 to MH05 are collected in a large industrial machine hall. The ground truth of relative pose is provided from the nonlinear least-squares batch solution over the Leica position and IMU measurements. The relative pose estimation of these sequences is challenging, because the industrial environment is unstructured and cluttered. Considering that the movement of the consecutive image pair is small, the image pairs for relative pose estimation are thinned out by an amount of one out of every four consecutive images. We also exclude the image pairs with insufficient motion in this experiment. The proposed solvers are compared with state-of-the-art solvers including 17pt-Li [35], 8pt-Kneip [24], and 6pt-Stewénius [46]. The SIFT detector [38] is used to establish the PCs of the image pair. All the solvers are integrated into the RANSAC framework to remove mismatches.
Table 11: Rotation and translation error on EuRoC sequences (unit: degree).
Seq. 17pt-Li [35] 8pt-Kneip [24] 6pt-Stew. [46]  6pt-Our-intra
MH01 0.136   3.055 0.156   3.214 0.187   4.181 0.130  2.961
MH02 0.129   2.806 0.132   2.796 0.198   4.193 0.127  2.579
MH03 0.199   2.422 0.187   2.517 0.236   3.789 0.181  2.376
MH04 0.195   3.159 0.178   3.237 0.229   5.440 0.193  3.105
MH05 0.186   3.124 0.163   2.940 0.241   4.464 0.158  2.892
Tab. 11 shows the rotation and translation error of the proposed 6pt-Our-intra solver for EuRoC sequences. The median error is used to evaluate the performance of solvers. It can be seen that the performance of the 6pt-Our-intra outperforms the comparative solvers 17pt-Li, 8pt-Kneip, and 6pt-Stewénius. It is shown that the proposed 6pt-Our-intra is well suited for the 6DOF relative pose estimation in the unmanned aerial vehicle environment.

### v2

@@ 0.F.3 Experiments on Synthetic Data
Figure 10: Rotation and translation error with varying image noise for the inter-camera case. We designed a simulated two-camera rig composed of two perspective cameras. The orientations of the two perspective cameras are roughly forward-facing with random perturbation. The resulting data is organized into three rows corresponding to forward, sideways, and random motions.
0.F.3.1 Synthetic Experiments for Two-camera Rigs
In this scenario, we design a simulated two-camera rig composed of two perspective cameras. The orientations of the two perspective cameras are roughly forward-facing with random perturbation. This setting is practical for autonomous driving with two front-facing cameras [10]. The performance of the solvers is tested in inter-camera and intra-camera configurations.
The baseline length between the two simulated cameras is fixed at $1$ meter. The multi-camera reference frame is defined at the center of the camera rig, with a translation distance of $3$ meters between two consecutive multi-camera reference frames. Each camera has a resolution of $640\times 480$ pixels, and a focal length of $400$ pixels. The principal points of the cameras are set at the image center. The synthetic scene consists of a ground plane and $50$ randomly generated planes. These planes are created within a cube-shaped region with dimensions of $[-5,5]\times[-5,5]\times[10,20]$ meters, defined in the corresponding axes of the multi-camera reference frame. Within the synthetic data, we randomly select $50$ PCs belonging to the ground plane, as well as a PC from each randomly generated plane. Therefore, a total of $100$ PCs are generated randomly in the synthetic scene.
The synthetic experiment consists of a series of $1000$ trials, wherein each trial entails the random generation of $100$ PCs. PCs are also randomly selected for the different solvers. The error evaluation is focused on determining the best relative pose, which yields the maximum inlier count within the RANSAC scheme. The RANSAC scheme also allows the selection of the most suitable candidate from multiple solutions. Rotation and translation errors are assessed using the median of errors. To induce different types of motion, the translation direction between two multi-camera reference frames is intentionally chosen to create forward, sideways, or random motions. For each motion, the second view is subjected to a random rotation, which is sequentially performed around three axes with the rotation angles ranging from $-10^{\circ}$ to $10^{\circ}$.
@@ 0.F.3.1 Synthetic Experiments for Two-camera Rigs
Figure 11: Rotation and translation error with varying image noise for the intra-camera case. We designed a simulated two-camera rig composed of two perspective cameras. The orientations of the two perspective cameras are roughly forward-facing with random perturbation. The resulting data is organized into three rows corresponding to forward, sideways, and random motions.
@@
Fig. 10 illustrates the performance comparison of different solvers with respect to image noise for the inter-camera case. A good method should have small median and variance for ${\varepsilon_{\mathbf{R}}}$, ${\varepsilon_{\mathbf{t}}}$, and $\varepsilon_{\mathbf{t},\text{dir}}$. The following observations are made: (1) The proposed 6pt-Our-inter solver demonstrates superior results compared to the comparative solvers. (2) The iterative optimization utilized in 8pt-Kneip is prone to convergence to local minima. This solver performs well in addressing forward and random motions. However, it does not perform well for the sideways motion. (3) The linear solver 17pt-Li with fewer calculations have less round-off error than the proposed 6pt-Our-inter solver in noise-free cases. However, our method has better accuracy than the linear solver with the influence of image noise. This is also consistent with real-world data experiments.
@@
Fig. 11 illustrates the performance comparison of different solvers with respect to image noise for the intra-camera case. The following observations are made: (1) Compared to the results in Fig. 10, the solvers using intra-camera correspondences generally perform worse than inter-camera correspondences, particularly in terms of recovering the metric scale of translation. (2) The proposed 6pt-Our-intra solver has better performance than the comparative solvers. (3) The 8pt-Kneip solver demonstrates satisfactory performance in the scenario of forward motion within multi-camera systems, but it displays diminished effectiveness in scenarios involving sideways and random motions.
0.F.3.2 Synthetic Experiments for Generalized Camera
@@
Table 10: The correspondence selection for different solvers. The pair $(i,i^{\prime})$ denotes a PC that is observed in the $i$-th camera in the first view and the $i^{\prime}$-th camera in the second view. The multiplier after a pair means the number of this match type.
@@
In this scenario, we design a simulated generalized camera composed of 12 omnidirectional cameras. The extrinsic parameters, including orientation and position, are totally random. Some of the comparison solvers use more PCs than the proposed solvers. The strategy of correspondence selection has a significant influence on their performance. We design a rule to select matches for different cases to guarantee fairness. Take the generic-camera configuration as an example. The match types are $(1,2)$, $(3,4)$, $(5,6)$, $(7,8)$, $(9,10)$, and $(11,12)$ for the proposed generic solver. Recall that the pair $(i,i^{\prime})$ denotes a PC that is observed in the $i$-th camera in the first view and the $i^{\prime}$-th camera in the second view. Solvers 17pt-Li, 8pt-Kneip, and 6pt-Stewénius cyclically build matches according to the order of $(1,2)$, $(3,4)$, $(5,6)$, $(7,8)$, $(9,10)$, and $(11,12)$. Tab. 10 shows the correspondence selection for different solvers.
0.F.4 Experiments on KITTI Dataset
Fig. 12 presents the empirical cumulative error distributions for sequence 00 in KITTI dataset. These distributions are derived from the same underlying values utilized to construct Table 2 in the paper. The results showcase that the proposed 6pt-Our-intra solver outperforms state-of-the-art solvers, displaying the most favorable overall performance. In addition, we have compared the suggested experiments with the solvers 6pt-Ventura [51] and 4pt-Sweeney [50], which utilize a prior for relative pose estimation. The median rotation and translation error on Sequence 00 of the KITTI dataset: 6pt-Ventura ($0.141^{\circ}$, $2.458^{\circ}$), 4pt-Sweeney ($0.056^{\circ}$, $2.289^{\circ}$). Our solver outperforms 6pt-Ventura in terms of accuracy. Since 4pt-Sweeney assumes known vertical direction as a prior, it achieves the highest accuracy. Given the proposed solvers do not exploit any motion prior, this comparative result does not actually harm our contribution.
@@ 0.F.4 Experiments on KITTI Dataset
0.F.5 Experiments on nuScenes Dataset
The performance evaluation of the solvers extends to the nuScenes dataset [6]. This dataset comprises consecutive keyframes captured from a multi-camera system consisting of six cameras, providing a comprehensive 360-degree field of view. For evaluation purposes, all keyframes from Part 1 are utilized, totaling 3,376 images. The ground truth is established through a lidar map-based localization scheme. Similar to the experiments conducted on the KITTI dataset, the SIFT detector [39] is employed to establish PCs between consecutive views in the six cameras. To deal with outlier matches in the feature correspondences, all the solvers are integrated into the RANSAC scheme.
Table 11: Rotation and translation error on nuScenes dataset (unit: degree).
@@ 0.F.5 Experiments on nuScenes Dataset
01 0.183  2.826 0.175   2.732 0.207   2.898 0.155  2.519
Tab. 11 illustrates the rotation and translation error of the 6pt-Our-intra solver on the Part1 of nuScenes dataset. We use median error to evaluate the performance of solvers. The experiment results highlight that the proposed 6pt-Our-intra solver achieves superior performance compared to other solvers. In addition, compared to experiments conducted on the KITTI dataset, this experiment provides further evidence supporting the direct applicability of the 6pt-Our-intra method for relative pose estimation in systems equipped with more cameras.
Appendix 0.G Limitations
The vanilla RANSAC is used due to its simplicity and few parameters. SOTA RANSAC variants are useful to push the envelope of performance for a method, but it is hard to guarantee fairness when comparing different solvers. Though the performance of the vanilla RANSAC is not as good as SOTA RANSAC variants, it is sufficient to provide a fair and clear benchmark for all minimal solvers.